OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (2, 8) and (x, 3)

OpenStudy (anonymous):

it can be 2, 16 5 1/3 or 12

OpenStudy (freckles):

well y varies inversely with x means for constant k we have: \[y=\frac{k}{x}\] first step is to plug in (2,8) to find k

OpenStudy (anonymous):

ok

OpenStudy (freckles):

then after you find k you can find x for when y is 3

OpenStudy (anonymous):

how would i find K?

OpenStudy (freckles):

by pluggin (2,8) as suggested above

OpenStudy (anonymous):

ok. so y= 2,8/k?

OpenStudy (freckles):

(2,8) tells us when x=2 that y=8

OpenStudy (anonymous):

ok so 8 = K/2

OpenStudy (freckles):

solve for k or i guess you are calling it K solve for K

OpenStudy (anonymous):

how do i find k?

OpenStudy (freckles):

by solving 8=k/2

OpenStudy (anonymous):

so k =4?

OpenStudy (freckles):

notice you can just multiply both sides by 2

OpenStudy (freckles):

hmm 4/2 =8 is not true

OpenStudy (anonymous):

so k is 2?

OpenStudy (freckles):

2/2=8 is not true

OpenStudy (anonymous):

16/2?

OpenStudy (freckles):

have you tried doing what I asked multiply both sides by 2

OpenStudy (freckles):

8=k/2 multiply both sides by 2 2(8)=k

OpenStudy (anonymous):

k=16

OpenStudy (freckles):

and as you see 16/2=8 is true yes k=16

OpenStudy (freckles):

\[y=\frac{16}{x}\]

OpenStudy (freckles):

now the question is to find x for when y=3

OpenStudy (anonymous):

so the answer is B?

OpenStudy (anonymous):

right?

OpenStudy (freckles):

have you found x for when y=3 yet?

OpenStudy (anonymous):

no...

OpenStudy (freckles):

well that is what the question ask you to do

OpenStudy (freckles):

it says to find x in the given ordered pair (x,3)

OpenStudy (freckles):

(x,3) says what is x for when y=3

OpenStudy (anonymous):

ok (2,3)

OpenStudy (freckles):

you already found the equation above

OpenStudy (freckles):

you just need to input the y is 3

OpenStudy (freckles):

and solve for x

OpenStudy (anonymous):

3= 16/8?

OpenStudy (freckles):

do you remember you found k was 16 so I replaced the k with 16 got you \[y=\frac{16}{x}\] re[place y with 3 and solve for x

OpenStudy (anonymous):

how do i solve X?

OpenStudy (freckles):

by solving the equation 3=16/x for x

OpenStudy (anonymous):

so that would be 5.3?

OpenStudy (freckles):

close in decimal form that would be 5.33333333333333333333 but I think they wanted you to write it in mix fraction form not decimal form

OpenStudy (anonymous):

what would that be as a mix fraction?

OpenStudy (freckles):

so you never wrote improper fractions as mix fractions before?

OpenStudy (freckles):

16/3 3 goes into 16, 5 times with a remainder of 1 \[\frac{16}{3}=5\frac{1}{3} \]

OpenStudy (anonymous):

OHH

OpenStudy (anonymous):

So the answer is D?

OpenStudy (freckles):

I don't know whatever choice is 5 and 1/3

OpenStudy (anonymous):

ok